OpenStudy (anonymous):

Explain, in complete sentences, the purpose of converting a radical expression into a rational exponent. Give an example

OpenStudy (jdoe0001):

do you know how to convert a radical to rational exponent?

OpenStudy (anonymous):

kind of

OpenStudy (jdoe0001):

ok.... do you know what say \(\bf x^n\times x^m = \square ?\)

OpenStudy (anonymous):

i cant see the exponents

OpenStudy (jdoe0001):

\( \huge x^n\times x^m = \square\)

OpenStudy (jdoe0001):

according to the exponents rules, what would that give you?

OpenStudy (anonymous):

im not sure tbh

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=kITJ6qH7jS0 <--- that covers those

OpenStudy (anonymous):

ok i appreciate it